OpenStudy (dtan5457):

Can anyone check a few questions for me? 1.If f(x)=-2x+7 and g(x)=x^2-2, f(g(3))=-7? 2. If f(x)=(2^x)-1 and g(x)=(x^2)-1, (f o g)(3)=63?

OpenStudy (anonymous):

First find f(g(x)) and then plug in 3.

OpenStudy (dtan5457):

the 2nd one?

OpenStudy (anonymous):

\[f(g(x)) = -2(x^2-2)+7\] now find f(g(3)) it's the same thing

Nnesha (nnesha):

and also you can find g(3) first and then plug that into f(g(x))

OpenStudy (anonymous):

Sure, that works to

OpenStudy (dtan5457):

I put g(3) first

OpenStudy (anonymous):

Basically it's telling you to take g(x) and plug it in f(x)

OpenStudy (dtan5457):

Then put that into f(x)

OpenStudy (dtan5457):

So I got g(3)=- f(g)(3)=-2(7)+7=-7

OpenStudy (dtan5457):

@Nnesha

Nnesha (nnesha):

loooks good :)

OpenStudy (dtan5457):

Now the 2nd one

OpenStudy (dtan5457):

How do I even plug in that one correctly?

Nnesha (nnesha):

oky for 2nd again first do g(3) replace x by 3 for g(x)

OpenStudy (dtan5457):

What happens if I first plug g into f?

OpenStudy (dtan5457):

(2^x-1)-1?

OpenStudy (dtan5457):

I mean (2^x^2-1)-1

Nnesha (nnesha):

okay back so for second one first you have to replace x by 3 into g(x)

Nnesha (nnesha):

that can be hard bec 2^x^2

Nnesha (nnesha):

idk if that's work or not but we did like this at school

OpenStudy (dtan5457):

So if it was put 3 into (g)x then, take that value to f(x) i get 3(g)x=8 right?

OpenStudy (dtan5457):

then 2^8-1=255?

OpenStudy (dtan5457):

if i put g(x) into f(x) first you get 2^(x^-1)-1

OpenStudy (dtan5457):

2^(x^2-1)-1

OpenStudy (dtan5457):

so same thing right?

Nnesha (nnesha):

a 3 power 2 is 9 and 9 -1 equal 8